OpenStudy (anonymous):

how do you divide 4 by 12

OpenStudy (nessapetrenko):

it's 3

OpenStudy (anonymous):

4 goes into 12 3 times. 4 + 4 +4 = 12 4x3 is 12. so 12/4 = 3

OpenStudy (anonymous):

no, you have to add a decimal point if you're doing long division. Is it 4 divided by 12 or 12 divided by 4? because 12 divided by 4 is 3

OpenStudy (anonymous):

if you're doing 4 divided by 12, it would be 0.33

OpenStudy (anonymous):

Yes it would be 0.33... :)

OpenStudy (anonymous):

Even if you do it on a calculator, 4 divided by 3 comes out 0.3333 repeating and you just round it :)

OpenStudy (anonymous):

Just leave it as a fraction or a repeating decimal. \(\frac{1}{3}\) or \(0.\overline{3}\).